# Magnetic Polarisability of Octet Baryons via Lattice QCD

### v2

@@ Abstract
Drawing on recent advances in lattice-QCD background-field techniques, the magnetic polarisability of octet baryons is calculated from the first principles of QCD. The results are presented in the context of new constituent quark-model calculations providing a framework for understanding the lattice results and a direct comparison with simulation results at unphysical quark masses. Using smeared quark sources, low-lying Laplacian eigenmode projection and final-state Landau mode projection, considerable attention is devoted to ensuring single-state isolation in the lattice correlation functions. We also introduce new weighting methods to reduce the sensitivity to correlation-function fits, averaging over many fits based on merit drawn from the full correlated $\chi^{2}$ of the fits. The techniques are implemented on the $32^{3}\times 64$, $2+1$-flavour dynamical-fermion lattices provided by the PACS-CS collaboration following the introduction of uniform magnetic fields quantised to the lowest nontrivial values available. After some fine tuning of the constituent quark model parameters, we find the model captures the patterns observed in the lattice QCD results very well, providing important insights into the physics underpinning the magnetic polarisabilities. Finally, comparison with the most recent results from experiment proceeds through an effective field theory formalism which incorporates estimates of finite-volume corrections and small electro-quenching corrections as the results are brought to the physical point. We find excellent agreement with experiment where available, including the proton and neutron polarisabilities.
@@ I Introduction
The magnetic polarisability is the second order response of an electrically charged composite particle to an external magnetic field. It encapsulates the manner in which the internal structure of the particle is changed by the field. From a perturbative point of view, the second-order process induces virtual electromagnetic transitions to nearby excitations in the hadron spectrum and probes the distribution of quarks within the hadron. In a nonperturbative sense, these virtual transitions act to change the structure of the hadron and thus the energy of the particle as it resides in the magnetic field.
@@
Our focus is on lattice QCD calculations of the magnetic polarisability of octet baryons. We draw on recent advances in lattice-QCD background-field techniques, established in an analysis of nucleon polarisabilities Bignell et al. (2020). In particular, smeared quark sources are utilised to capture the QCD aspects of the hadron structure. Then low-lying QCD+QED SU(3)$\times$×U(1) Laplacian eigenmode projection is considered at the quark level to capture the low-energy response of individual electrically-charged quarks to QCD and the external magnetic field. Electrically charged baryons will be in a superposition of Landau-level states and therefore final-state colour-singlet U(1) Landau-mode projection is used at the sink to ensure the baryon is in the lowest-lying Landau level Bignell et al. (2018, 2020); Tiburzi and Vayl (2013). We extend the previous analysis by devoting considerable attention to ensuring single-state isolation in the lattice correlation functions. We further introduce new weighting methods to reduce the sensitivity to correlation-function fits, averaging over many fits based on merit determined by the full covariance-matrix based $\chi^{2}$ of the fits.
@@
To help in understanding the pattern of octet-baryon magnetic polarisabilities observed, we consider the simple nonrelativistic constituent quark model and extend calculations of the nucleon polarisabilities Capstick and Keister (1992); Lipkin and Moinester (1992); Bhaduri and Warke (1988) to include the hyperons of the baryon octet. As we will demonstrate, chiral nonanalytic behaviour is subtle at the quark masses we consider on the lattice. Therefore we anticipate this simple model will be of utility in understanding the competing effects that generate some complexity in the pattern of magnetic polarisabilities observed in lattice QCD. Qualitatively, it explains the large difference between negatively charged baryons and the remainder of the octet. After some fine tuning of the constituent quark model parameters, we find the model captures the patterns observed in the lattice QCD results very well, providing important insights into the physics underpinning the magnetic polarisabilities.
@@
Our approach to the magnetic polarisability in lattice QCD centres on the uniform background magnetic field method. Historically, this approach has been the method of choice for determining baryon magnetic polarisabilities Zhou et al. (2003); Lee et al. (2006); Primer et al. (2014); Chang et al. (2015); Bignell et al. (2018, 2020). As a leading effect in the expansion of the baryon energy in terms of the magnetic field strength, the background field must be weak to ensure higher ${\cal O}(B^{3})$ terms are small. Thus, subtle shifts in the energy of the baryon induced by the magnetic field need to be related to the polarisability. To reveal the effects of the magnetic field, correlated ratios are constructed to enable QCD based fluctuations to cancel.
@@
Our techniques are implemented on the $32^{3}\times 64$, $2+1$-flavour dynamical-fermion lattices provided by the PACS-CS collaboration. QCD correlations are maintained by applying the external magnetic field to these existing gauge field configurations. A consequence of this is that our gauge fields are electro-quenched in that the sea-quarks are blind to the magnetic field. At the quark masses considered herein, effective field theory indicates the approximation is reasonable, and we use effective field theory to estimate the small electro-quenching corrections Deshmukh and Tiburzi (2018); Hall et al. (2014); Bignell et al. (2020). We also draw on effective field theory to estimate the small finite-volume corrections to our lattice QCD calculations Hall et al. (2014); Bignell et al. (2020).
@@
In constructing the required correlated ratios, several two-point correlation functions are combined. These include the zero-field correlator and spin-field aligned and anti-aligned baryon correlators at finite magnetic-field strength. These all display different behaviour as a function of Euclidean time and their combination can hide excited-state contamination effects. To address this, we examine the Euclidean-time behaviour of each of these underlying correlators to ensure that the effective energy of each two-point correlator has reached single-state isolation through Euclidean time evolution. We find that in many cases plateau-like behaviour in the correlated ratios precedes the single-state isolation of the individual underlying correlation functions. Thus the results presented herein provide an important step forward in suppressing contamination from excited states.
@@
We also carefully examine the dependence of the results on the selection of the fit window where a significant dependence can arise in some cases. To address this we introduce a weighting method to reduce the sensitivity to fit-window selection. Following the ideas presented in Ref. Beane et al. (2021), we average over many fit windows with their weight based on merit calculated with the full covariance-matrix based $\chi^{2}$ of the fits.
@@
We find the processes described above bring a new level of rigour to the polarisability results, producing the highest precision lattice QCD calculations of baryon magnetic polarisabilities to date.
@@
The presence of a magnetic field and the difference in the $u$ and $d$ quark charges breaks the isospin symmetry of equal mass quarks. The $\Lambda$ and $\Sigma^{0}$ baryons will mix in the magnetic field complicating the interpretation of the mass shifts in the background field approach considered herein. Correlation matrix techniques can be introduced to address the state isolation issue. However, together these issues add additional layers of complexity that is not required for the other six members of the baryon octet. These are the so-called outer members of the octet, composed with a doubly-represented quark flavour and a singly-represented flavour. It is these baryons that hold the focus of the present investigation.
@@
Recently four-point function methods have been considered in the calculation of the magnetic polarisability Wilcox and Lee (2021); Lee et al. (2023). This time the baryons are probed perturbatively by two electromagnetic-current insertions. As such, the baryon states are eigenstates of QCD alone and the $\Lambda$ and $\Sigma^{0}$ baryons do not mix. Thus, this alternative approach presents a significant advantage to understanding the magnetic polarisabilities of the $\Lambda$ and $\Sigma^{0}$ baryons and we anticipate future calculations for these baryons.
@@
The presentation of this work is as follows. In Sec. II we derive the generalised expression for the magnetic polarisability of an octet baryon in a simple constituent quark model. Highlighting the competing effects of magnetic transitions and the distribution of charge within the baryon, we derive predictions for the magnetic polarisabilities of the outer octet baryons and use these to provide context for the results of the lattice QCD calculations. In Sec. III we describe our implementation of the background field for the lattice QCD calculation, emphasising the co-location of the field and baryon correlator origin. In Sec. IV we discuss the extraction of the magnetic polarisability from two-point baryon correlation functions and ratios. In Sec. V we discuss relevant details of the lattice simulation including the ensembles used, the quark level SU(3)$\times$×U(1) projection at the quark-propagator sink, and the U(1) hadronic projection which allows for proper handling of Landau-level physics. In Sec. VI we discuss our advances in the correlation-function fitting methods, including the examination of the excited state contamination in the underlying correlation functions and the weighted averaging of fit windows. In Sec. VII we present the results of the lattice calculation and utilise them to improve the quark model such that it can capture the essential physics governing baryon magnetic polarisabilities. In Sec. VIII we use chiral effective field theory to incorporate electro-quenching corrections and estimate finite-volume corrections before extrapolating the lattice results to the physical regime. In Sec. IX we compare our findings with experimental measurements and previous determinations of octet-baryon magnetic polarisabilities. In Sec. X we briefly examine the proton-neutron magnetic polarisability difference. In Sec. XI we summarise our findings.
@@ II Constituent Quark Model
The constituent quark model is renowned for its capacity to provide a simple explanation for the pattern of baryon magnetic moments observed in experiment. Founded on SU(6) spin-flavour symmetry, symmetric ground-state wave functions with zero angular momentum, and anti-symmetry via SU(3) colour-singlet states, the model captures the essence of the physics with subtle corrections arising from the environment sensitivity of the constituent quark moments Leinweber et al. (1991); Boinepalli et al. (2006). We consider the constituent quark model predictions for octet-baryon magnetic polarisabilities in a similar spirit, hoping to capture the predominant features in a simple model and providing a framework for the description of more subtle corrections.
II.1 Magnetic polarisabilities from the quark model
Bhaduri et al. Bhaduri and Warke (1988) present an expression for the magnetic polarisability of the proton in the constituent quark model. Here we review the derivation of this expression and generalise it for an arbitrary octet baryon. In doing so, we aim to provide insight into the salient features of QCD that give rise to the magnetic polarisabilities of ground-state baryons, at quark masses where pion dressings of the baryons are of secondary importance Hall et al. (2014); Bignell et al. (2020).
A quark model expression for the magnetic polarisability may be obtained using Rayleigh-Schrödinger perturbation theory, reviewed in Ref. Sakurai and Commins (1995) for example. We consider an unperturbed Hamiltonian $\mathcal{H}_{0}$, and label the energy eigenstates of this Hamiltonian |$\ket{n_{i}}$
@@ II.1 Magnetic polarisabilities from the quark model
$(\mathcal{H}_{0}+\lambda\mathcal{H}_{\rm int})\ket{k_{i}}=E^{\prime}_{i}\ket{k_{i}},\quad i=1,2,3,\ldots\,.$ ((2))
@@
The magnetic polarisability is associated with terms quadratic in the magnetic field strength and thus $e^{2}$ and we’ll see both correction terms of Eq. (3) provide a contribution. Focusing on terms proportional to $e^{2}$, we define the magnetic polarisability, $\beta$, via
@@
Here we are purely interested in ground state baryons. As such, the quarks are in a relative S-wave state and have zero angular momentum causing that term to vanish, leaving
@@
where we now see that the $e^{2}$ term contains the position operators. Thus the magnetic polarisability contains contributions probing the distribution of quarks within the hadron.
@@
This interaction Hamiltonian is then substituted into Eq. (3). Such a substitution requires the determination of two matrix elements. Firstly
@@
where $\mu_{\mathcal{B}}$ is the magnetic moment of the baryon, and $\mu_{D}$ and $\mu_{S}$ are the intrinsic magnetic moments of the doubly-represented and singly-represented quark flavours respectively. In the weak field limit, spherical symmetry of the $S$-wave states provides $\expectationvalue{x^{2}}=\expectationvalue{y^{2}}=\frac{1}{3}\expectationvalue{r^{2}}$.
The second matrix element of interest in Eq. (3) contains an exited state of the octet baryon in the ket. The sum is estimated by considering dominance of the first nearby excitation, the corresponding decuplet baryon state. This matrix element generates terms proportional to $e^{2}$ via
@@
where $D$ and $S$ once again label the constituent quark flavours for the doubly- and singly-represented sectors; for example for the proton, $\mu_{D}-\mu_{S}=\mu_{u}-\mu_{d}$.
@@
where the definition of Eq. (5) has been used. Thus, the constituent quark model prediction for the magnetic polarisability of octet baryon $\mathcal{B}$ composed of three quark flavours denoted by $f$ is
@@
where $\alpha={e^{2}}/{4\pi}$ is the fine structure constant. The magnetic polarisability has its origin in two competing terms which we have defined as $\beta_{1}$, a magnetic transition term, and $\beta_{2}$ probing the distribution of quarks within the baryon. The two terms and their opposing signs in this expression highlight the complex nature of the baryon polarisability. This is the constituent quark model description for the magnetic polarisability of an octet baryon with two quarks of one flavour and another quark with a different flavour.

### v3

# Magnetic polarizability of octet baryons via lattice QCD

@@ Abstract
Drawing on recent advances in lattice-QCD background-field techniques, the magnetic polarizability of octet baryons is calculated from the first principles of QCD. The results are presented in the context of new constituent quark-model calculations providing a framework for understanding the lattice results and a direct comparison with simulation results at unphysical quark masses. Using smeared quark sources, low-lying Laplacian eigenmode projection and final-state Landau-mode projection, considerable attention is devoted to ensuring single-state isolation in the lattice correlation functions. We also introduce new weighting methods to reduce the sensitivity to correlation-function fits, averaging over many fits based on merit drawn from the full correlated $\chi^{2}$ of the fits. The techniques are implemented on the $32^{3}\times 64$, $2+1$-flavor dynamical-fermion lattices provided by the PACS-CS Collaboration following the introduction of uniform magnetic fields quantized to the lowest nontrivial values available. After some scaling of the constituent quark model parameters, we find the model captures the patterns observed in the lattice QCD results very well, providing important insights into the physics underpinning the magnetic polarizabilities. Finally, comparison with the most recent results from experiment proceeds through an effective field theory formalism which incorporates estimates of finite-volume corrections and small electroquenching corrections as the results are brought to the physical point. We find excellent agreement with experiment where available, including the proton and neutron polarizabilities.
@@ I Introduction
The magnetic polarizability is the second-order response of an electrically charged composite particle to an external magnetic field. It encapsulates the manner in which the internal structure of the particle is changed by the field. From a perturbative point of view, the second-order process induces virtual electromagnetic transitions to nearby excitations in the hadron spectrum and probes the distribution of quarks within the hadron. In a nonperturbative sense, these virtual transitions act to change the structure of the hadron and thus the energy of the particle as it resides in the magnetic field.
@@
Our focus is on lattice QCD calculations Gross et al. (2023) of the magnetic polarizability of octet baryons. We draw on recent advances in lattice-QCD background-field techniques, established in an analysis of nucleon polarizabilities Bignell et al. (2020). In particular, smeared quark sources are utilized to capture the QCD aspects of the hadron structure. Then low-lying QCD+QED SU(3)$\times$×U(1) Laplacian eigenmode projection is considered at the quark level to capture the low-energy response of individual electrically charged quarks to QCD and the external magnetic field. Electrically charged baryons will be in a superposition of Landau-level states and therefore final-state color-singlet U(1) Landau-mode projection is used at the sink to ensure the baryon is in the lowest-lying Landau level Bignell et al. (2018, 2020); Tiburzi and Vayl (2013). We extend the previous analysis by devoting considerable attention to ensuring single-state isolation in the lattice correlation functions. We further introduce new weighting methods to reduce the sensitivity to correlation-function fits, averaging over many fits based on merit determined by the full covariance-matrix-based $\chi^{2}$ of the fits.
@@
To help in understanding the pattern of octet-baryon magnetic polarizabilities observed, we consider the simple nonrelativistic constituent quark model and extend calculations of the nucleon polarizabilities Capstick and Keister (1992); Lipkin and Moinester (1992); Bhaduri and Warke (1988) to include the hyperons of the baryon octet. As we will demonstrate, chiral nonanalytic behavior is subtle at the quark masses we consider on the lattice. Therefore, we anticipate this simple model will be of utility in understanding the competing effects that generate some complexity in the pattern of magnetic polarizabilities observed in lattice QCD. Qualitatively, it explains the large difference between negatively charged baryons and the remainder of the octet. After some scaling of the constituent quark model parameters, we find the model captures the patterns observed in the lattice QCD results very well, providing important insights into the physics underpinning the magnetic polarizabilities.
@@
Our approach to the magnetic polarizability in lattice QCD centers on the uniform background magnetic-field method. Historically, this approach has been the method of choice for determining baryon magnetic polarizabilities Zhou et al. (2003); Lee et al. (2006); Primer et al. (2014); Chang et al. (2015); Bignell et al. (2018, 2020). As a leading effect in the expansion of the baryon energy in terms of the magnetic-field strength, the background field must be weak to ensure higher ${\cal O}(B^{3})$ terms are small. Thus, subtle shifts in the energy of the baryon induced by the magnetic field need to be related to the polarizability. To reveal the effects of the magnetic field, correlated ratios are constructed to enable QCD-based fluctuations to cancel.
@@
Our techniques are implemented on the $32^{3}\times 64$, $2+1$-flavor dynamical-fermion lattices provided by the PACS-CS Collaboration. QCD correlations are maintained by applying the external magnetic field to these existing gauge field-configurations. A consequence of this is that our gauge fields are ”electroquenched” in that the sea-quarks are blind to the magnetic field. At the quark masses considered herein, effective field theory indicates the approximation is reasonable, and we use effective field theory to estimate the small electroquenching corrections Deshmukh and Tiburzi (2018); Hall et al. (2014); Bignell et al. (2020). We also draw on effective field theory to estimate the small finite-volume corrections to our lattice QCD calculations Hall et al. (2014); Bignell et al. (2020).
@@
In constructing the required correlated ratios, several two-point correlation functions are combined. These include the zero-field correlator and spin-field aligned and antialigned baryon correlators at finite magnetic-field strength. These all display different behavior as a function of Euclidean time and their combination can hide excited-state contamination effects. To address this, we examine the Euclidean time behavior of each of these underlying correlators to ensure that the effective energy of each two-point correlator has reached single-state isolation through Euclidean time evolution. We find that in many cases plateaulike behavior in the correlated ratios precedes the single-state isolation of the individual underlying correlation functions. Thus the results presented herein provide an important step forward in suppressing contamination from excited states.
@@
We also carefully examine the dependence of the results on the selection of the fit window where a significant dependence can arise in some cases. To address this we introduce a weighting method to reduce the sensitivity to fit window selection. Following the ideas presented in Ref. Beane et al. (2021), we average over many fit windows with their weight based on merit calculated with the full covariance-matrix-based $\chi^{2}$ of the fits.
@@
We find the processes described above bring a new level of rigor to the polarizability results, producing the highest precision lattice QCD calculations of baryon magnetic polarizabilities to date.
@@
The presence of a magnetic field and the difference in the $u$ and $d$ quark charges breaks the isospin symmetry of equal mass quarks. The $\Lambda$ and $\Sigma^{0}$ baryons will mix in the magnetic field complicating the interpretation of the mass shifts in the background-field approach considered herein. Correlation matrix techniques can be introduced to address the state isolation issue. However, together these issues add additional layers of complexity that are not required for the other six members of the baryon octet. These are the so-called outer members of the octet, composed with a doubly represented quark flavor and a singly represented flavor. It is these baryons that hold the focus of the present investigation.
@@
Recently four-point function methods have been considered in the calculation of the magnetic polarizability Wilcox and Lee (2021); Lee et al. (2023). This time the baryons are probed perturbatively by two electromagnetic-current insertions. As such, the baryon states are eigenstates of QCD alone and the $\Lambda$ and $\Sigma^{0}$ baryons do not mix. Thus, this alternative approach presents a significant advantage to understanding the magnetic polarizabilities of the $\Lambda$ and $\Sigma^{0}$ baryons and we anticipate future calculations for these baryons.
@@
The presentation of this work is as follows. In Sec. II we derive the generalized expression for the magnetic polarizability of an octet baryon in a simple constituent quark model. Highlighting the competing effects of magnetic transitions and the distribution of charge within the baryon, we derive predictions for the magnetic polarizabilities of the outer octet baryons and use these to provide context for the results of the lattice QCD calculations. In Sec. III we describe our implementation of the background field for the lattice QCD calculation, emphasizing the colocation of the field and baryon correlator origin. In Sec. IV we discuss the extraction of the magnetic polarizability from two-point baryon correlation functions and ratios. In Sec. V we discuss relevant details of the lattice simulation including the ensembles used, the quark level SU(3)$\times$×U(1) projection at the quark propagator sink, and the U(1) hadronic projection which allows for proper handling of Landau-level physics. In Sec. VI we discuss our advances in the correlation-function fitting methods, including the examination of the excited-state contamination in the underlying correlation functions and the weighted averaging of fit windows. In Sec. VII we present the results of the lattice calculation and utilize them to improve the quark model such that it can capture the essential physics governing baryon magnetic polarizabilities. In Sec. VIII we use chiral effective field theory to incorporate electroquenching corrections and estimate finite-volume corrections before extrapolating the lattice results to the physical regime. In Sec. IX we compare our findings with experimental measurements and previous determinations of octet-baryon magnetic polarizabilities. In Sec. X we briefly examine the proton-neutron magnetic polarizability difference. In Sec. XI we summarize our findings.
@@ II Constituent Quark Model
The constituent quark model is renowned for its capacity to provide a simple explanation for the pattern of baryon magnetic moments observed in experiment. Founded on SU(6) spin-flavor symmetry, symmetric ground-state wave functions with zero angular momentum, and antisymmetry via SU(3) color-singlet states Close (1979), the model captures the essence of the physics with subtle corrections arising from the environment sensitivity of the constituent quark moments Leinweber et al. (1991); Boinepalli et al. (2006). We consider the constituent quark model predictions for octet-baryon magnetic polarizabilities in a similar spirit, hoping to capture the predominant features in a simple model and providing a framework for the description of more subtle corrections.
II.1 Magnetic polarizabilities from the quark model
Bhaduri et al. Bhaduri and Warke (1988) present an expression for the magnetic polarizability of the proton in the constituent quark model. Here we review the derivation of this expression and generalize it for an arbitrary octet baryon. In doing so, we aim to provide insight into the salient features of QCD that give rise to the magnetic polarizabilities of ground-state baryons, at quark masses where pion dressings of the baryons are of secondary importance Hall et al. (2014); Bignell et al. (2020).
A quark model expression for the magnetic polarizability may be obtained using Rayleigh-Schrödinger perturbation theory, reviewed in Ref. Sakurai and Napolitano (2020) for example. We consider an unperturbed Hamiltonian $\mathcal{H}_{0}$, and label the energy eigenstates of this Hamiltonian |$\ket{n_{i}}$
@@ II.1 Magnetic polarizabilities from the quark model
$(\mathcal{H}_{0}+\lambda\mathcal{H}_{\rm int})\ket{k_{i}}=E^{\prime}_{i}\ket{k_{i}},\quad i=1,2,3,\ldots\,,$ ((2))
@@
The magnetic polarizability is associated with terms quadratic in the magnetic-field strength and thus $e^{2}$ and we will see both correction terms of Eq. (3) provide a contribution. Focusing on terms proportional to $e^{2}$, we define the magnetic polarizability, $\beta$ via
@@
Here we are purely interested in ground-state baryons. As such, the quarks are in a relative $S$-wave state and have zero angular momentum causing that term to vanish, leaving
@@
where we now see that the $e^{2}$ term contains the position operators. Thus, the magnetic polarizability contains contributions probing the distribution of quarks within the hadron.
@@
This interaction Hamiltonian is then substituted into Eq. (3). Such a substitution requires the determination of two matrix elements. First,
@@
where $\mu_{\mathcal{B}}$ is the magnetic moment of the baryon, and $\mu_{D}$ and $\mu_{S}$ are the intrinsic magnetic moments of the doubly represented and singly represented quark flavors respectively. In the weak-field limit, spherical symmetry of the $S$-wave states provides $\expectationvalue{x^{2}}=\expectationvalue{y^{2}}=\frac{1}{3}\expectationvalue{r^{2}}$.
The second matrix element of interest in Eq. (3) contains an excited state of the octet baryon in the ket. The sum is estimated by considering dominance of the first nearby excitation, the corresponding decuplet baryon state. This matrix element generates terms proportional to $e^{2}$ via
@@
where $D$ and $S$ once again label the constituent quark flavors for the doubly and singly represented sectors; for example for the proton, $\mu_{D}-\mu_{S}=\mu_{u}-\mu_{d}$.
@@
where the definition of Eq. (5) has been used. Thus, the constituent quark model prediction for the magnetic polarizability of octet baryon $\mathcal{B}$ composed of three quark flavors denoted by $f$ is
@@
where $\alpha={e^{2}}/{4\pi}$ is the fine structure constant. The magnetic polarizability has its origin in two competing terms which we have defined as $\beta_{1}$, a magnetic transition term, and $\beta_{2}$ probing the distribution of quarks within the baryon. The two terms and their opposing signs in this expression highlight the complex nature of the baryon polarizability. This is the constituent quark model description for the magnetic polarizability of an octet baryon with two quarks of one flavor and another quark with a different flavor.